The weights of students in a certain school are normally distributed with a mean weight of 66 kg. 10% have a weight greater than 70kg. What percentage of students weighs between 52kg and 66kg?   


1
Expert's answer
2022-04-11T18:12:34-0400

"P(X>70)=0,1;\\
P(X<70)=1-0.1=0.9."

The closest to 0.9 value in the z-table is 0.8997 for z=1.28.

"z=\cfrac{x-\mu}{\sigma},\\
\sigma=\cfrac{x-\mu}{z}=\cfrac{70-66}{1.28}=3.125;"

"z_1=\cfrac{52-66}{3.125}=-1.92;\\
z_2=\cfrac{66-66}{3.125}=0;\\
P(52<X<66)=P(-1.92<Z<0)=\\
=P(Z<0)-P(Z<-1.92)=\\
=0.5000-0.0274=0.4726 \text{ (from z-table).}"